OpenStudy (anonymous):

What is the solution to the following equation? (x - 3)2 = 14

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

I guess this equation should be \[(x-3)^2=14\] \[x^2-6x+9=14\] \[x^2-6x+9-14=0\] \[x^2-6x-5=0\] With Bhaskara \[\boxed{\boxed{x_1=3+\sqrt{14}~~and~~x_2=3-\sqrt{14}}}\]

OpenStudy (idkwut):

Ah, opps. Dexter may be right on this one. Didn't realize that the 2 was an exponent. Sorry.